# Adaptive Parameter-Free Robust Learning using Latent Bernoulli Variables

### v1

@@ Abstract
We present an efficient parameter-free approach for statistical learning from corrupted training sets. We identify corrupted and non-corrupted samples using latent Bernoulli variables, and therefore formulate the robust learning problem as maximization of the likelihood where latent variables are marginalized out. The resulting optimization problem is solved via variational inference using an efficient Expectation-Maximization based method. The proposed approach improves over the state-of-the-art by automatically inferring the corruption level and identifying outliers, while adding minimal computational overhead. We demonstrate our robust learning method on a wide variety of machine learning tasks including online learning and deep learning where it exhibits ability to adapt to different levels of noise and attain high prediction accuracy.
44footnotetext: aleksandr.karakulev@it.uu.se11footnotetext: Co-senior authors
@@ 1 Introduction
Several statistical learning problems are formulated as estimation of parameters $\bm{\theta}\in\Theta$ of a probabilistic model by maximizing its likelihood function $p(\bm{z}|\bm{\theta})$ for observation $\bm{z}\sim p(\bm{z})$. It is more convenient to define the loss function as the negative log-likelihood $\ell_{\bm{\theta}}(\bm{z})=-\ln p(\bm{z}|\bm{\theta})$, and solve the equivalent minimization problem. Hence, given a training set of $n$ independent observations $Z=\{\bm{z}_{i}\}_{i=1}^{n}$, the learning problem consists of finding the solution,
@@
However, real-world data is often corrupted – samples arise from a mixture of the true distribution $p(\bm{z})$ and a corrupting source $q(\bm{z})$,
$\widetilde{p}(\bm{z})=(1-\varepsilon)p(\bm{z})+\varepsilon q(\bm{z}),$ (2)
@@
leading to a suboptimal solution $\bm{\theta}_{\text{ML}}$. In this contaminated mixture, outliers $\bm{z}\sim q(\bm{z})$ may result from inaccurate measurements, errors or oversight in data acquisition or labeling, or even malicious attacks. The situation is further complicated by the fact that the corrupting distribution $q(\bm{z})$ is typically unknown, and the level of corruption $\varepsilon$, difficult to determine. Equation (2) used herein follows the well-known Huber contamination model from the robust statistics literature [10, 15]. Existing approaches require an estimation of the noise structure or the corruption level $\varepsilon$. This imposes restrictions on using these methods in settings where such pre-processing becomes impractical, e.g., in online learning, wherein the assumption about $\varepsilon$ needs to be optimized in continuously arriving data. In Figure 1, we show an example: the accuracy of binary classification trained on data that is continuously collected from the HAR dataset [8], and subsequently corrupted with varying number of randomly flipped labels in each batch (details in Section 4.2).
@@
Figure 1: Classification of online streaming data with varying number of outliers. The parameter-free nature of our approach (RLVI) allows for automatic identification of outliers when learning from batches of data with different $\varepsilon$. Our method is robust and thus has higher accuracy than the standard stochastic optimization of the likelihood (SGD).
Contribution. We present a principled parameter-free robust learning approach for settings involving likelihood maximization that introduces minimal computational overhead, and is general in applicability with respect to model or problem type. We demonstrate the approach in online learning and deep learning settings, where it meets or exceeds state-of-the-art results on benchmark datasets.
Related work. While some approaches address learning from corrupted data with a special loss function, e.g., the Huber loss [9], others construct certain criteria to separate samples from the true distribution, from the corrupted ones [1, 2]. Recently, new general approaches, such as Sever [5] and Robust Risk Minimization (RRM) [18], were introduced. Sever is a general gradient-based learning method that penalizes the gradient components of the loss function corresponding to outliers. RRM involves obtaining sample weights from the constraint on the entropy of a re-weighted empirical distribution, and minimizing the corresponding modified empirical risk. Both methods are not model-specific and do not require the exact value of $\varepsilon$, rather only an upper-bound. Robust Risk Minimization does not require calibrating multiple hyperparameters like Sever, but requires assessing the level of corruption by estimating its maximum value.
In the recent past, learning from corrupted data has also received attention in the field of deep learning, and several approaches have been proposed to train a neural network in a robust way. For instance, one can use the correction of the regular loss function [24, 19]. Alternatively, some approaches directly estimate the noise transition matrix [6, 19]. Others utilize an additional neural network to identify and prevent overfitting of the initial model, as in [11], and to obtain a better accuracy by exchanging the information between two simultaneously trained networks, as in [7] and [20]. Another option is to combine techniques that are specific to deep learning – early-stopping [22] or dropout [23], with a criterion that aims to eliminate corrupted samples. However, training a neural network in the regular context where the contamination issue is not addressed, already requires certain hyperparameters to improve generalization, e.g., learning rate, batch-size, number of layers, etc. The aforementioned approaches introduce additional parameters to combat the problem of corrupted labels, including $\varepsilon$, which makes their performance dependent on the efficacy of these parameters and the underlying assumptions.
@@
Robust Learning via Variational Inference. The proposed approach, denoted as RLVI, employs latent Bernoulli variables to identify corrupted and non-corrupted training data samples. Distribution of the latent variables is characterized from training data by maximizing the variational lower bound of marginal likelihood (variational inference) [16]. This allows detection of outliers and computation of the optimal corruption level $\varepsilon$ automatically, subsequently learning the model $\bm{\theta}$ in a robust way. The problem formulation entailing RLVI is theoretically simple, and also leads to efficient and straightforward implementation.
The paper is structured as follows. Section 2 presents the problem formulation – introducing the latent variables enabling the marginal likelihood approach. Section 3 describes the Expectation-Maximization based method, realizing variational inference, and its extension to stochastic optimization for online learning and deep learning problems. Section 4 validates the efficacy of the proposed approach against existing approaches on standard statistical learning problems (linear regression and classification, PCA, online learning) [18, 10, 5] and image classification [7, 22, 20, 23].
@@ 2 Problem Formulation
Central to our approach is the introduction of a latent variable $t_{i}$ for each observation $\bm{z}_{i}\in Z$ such that,
@@
Knowing the values of these variables would allow us to avoid minimizing the losses on outliers in the dataset in (1) by simply dropping the corresponding terms from the sum. In other words, latent variables enable us to define a likelihood function with respect to the non-corrupted data, such that,
@@
where $\bm{t}=[t_{1},t_{2},\dots,t_{n}]$. Optimization with respect to this likelihood function implies a combinatorial search over the latent variables $\bm{t}$, which is computationally infeasible even for moderate $n$.
@@
$\bm{\theta}_{\star}=\underset{\bm{\theta}}{\arg\max}\,\max_{\varepsilon}p(Z|\bm{\theta},\varepsilon),$ (7)
addresses the problem of robust learning of $\bm{\theta}$ while obviating the need for specifying $\varepsilon$ or a combinatorial search over $\bm{t}$. We now turn to developing a method to solve (7) in an efficient manner.
@@ 3 Method
We begin by introducing a variational lower bound on the marginal likelihood in (7) and then turn to implementation-specific aspects of the derived method.
Variational inference. To solve the formulated optimization problem, we need a tractable way to compute the sum over $\bm{t}$ in (6). Using Bayes’ rule, we express the marginal likelihood as,
@@
where the denominator is the posterior distribution of latent variables $\bm{t}$ given the data $Z$, specified model $\bm{\theta}$, and the corruption level $\varepsilon$. As this posterior is intractable, we consider its variational approximation denoted $r(\bm{t}|\bm{\pi})$. Specifically, we use a distribution of $n$ independent Bernoulli variables with probabilities $\bm{\pi}=[\pi_{1},\dots,\pi_{n}]$,
@@
Consequently, instead of maximizing the marginal likelihood function $p(Z|\bm{\theta},\varepsilon)$ directly, we apply the variational inference framework, and optimize the so-called evidence lower-bound (ELBO) [16]. This lower bound has the following general form,
$\ln{p(Z|\bm{\theta},\varepsilon)}\geq\mathbb{E}_{r(\bm{t}|\bm{\pi})}\Big{[}\ln{p(Z|\bm{\theta},\bm{t})}\Big{]}-\text{KL}\Big{[}r(\bm{t}|\bm{\pi})\big{|}\big{|}p(\bm{t}|\varepsilon)\Big{]},$ (10)
@@
Note that corruption level does not appear in the first term of the ELBO, and we can optimize the KL-divergence w.r.t. $\varepsilon$ directly and independently of the model $\bm{\theta}$. This results in ${n(1-\varepsilon)=\sum_{i}\pi_{i}}$. In other words, parameters $\pi_{i}$ defining the probability of each sample being non-corrupted, should sum up to the number of non-corrupted samples. We can therefore express the optimal corruption fraction as,
$\varepsilon=1-\langle\bm{\pi}\rangle,$ (13)
where $\langle\bm{\pi}\rangle$ is the average of the Bernoulli probabilities. Hence, we arrive at the following objective – the negative ELBO expressed as,
$\mathcal{L}(\bm{\theta},\bm{\pi})=\sum\limits_{i=1}^{n}\pi_{i}\ell_{\bm{\theta}}(\bm{z}_{i})+\pi_{i}\ln\frac{\pi_{i}}{\langle\bm{\pi}\rangle}+(1-\pi_{i})\ln\frac{1-\pi_{i}}{1-\langle\bm{\pi}\rangle}.$ (14)
In conclusion, the objective in (7) is replaced by the optimum of its variational bound, i.e.,
$\boxed{\bm{\theta}_{\text{VI}}=\underset{\bm{\theta}\in\Theta}{\arg\min}\,\min_{\bm{\pi}\in[0;1]^{n}}{\mathcal{L}(\bm{\theta},\bm{\pi})}.}$ (15)
Numerical optimization. The resulting optimization problem is solved using the block-wise Algorithm 1, which follows the general Expectation-Maximization (EM) scheme [3]. The E-step consists of minimizing $\mathcal{L}(\bm{\theta},\bm{\pi})$ in $\bm{\pi}$ for fixed parameters $\bm{\theta}$. In the M-step, the inferred probabilities are used to maximize the re-scaled log-likelihood, $-\sum_{i=1}^{n}\pi_{i}\ell_{\bm{\theta}}(\bm{z}_{i})$. As an example, in case of linear regression the latter step solves a weighted least-squares problem. For a classification task, we minimize the cross-entropy loss corrected with sample weights.
Note that the E-step can be performed efficiently, as the objective $\mathcal{L}(\bm{\theta},\bm{\pi})$ is convex in $\bm{\pi}$ (proof is in the Appendix). Therefore, to find the optimal parameters $\bm{\pi}$ for a fixed model, the derivative of $\mathcal{L}(\bm{\theta},\bm{\pi})$ is equated to zero w.r.t. $\pi_{j}$ for all $j=1,\dots,n$,
@@
This is a system of nonlinear equations for $\bm{\pi}$, which we solve with the fixed-point iteration procedure,
$\pi_{j}^{\text{new}}=\left(1+\dfrac{1-\langle\bm{\pi}^{\text{old}}\rangle}{\langle\bm{\pi}^{\text{old}}\rangle}e^{\ell_{\bm{\theta}}(\bm{z}_{j})}\right)^{-1}.$ (17)
Here, for the mean weight, we use the value from the previous fixed-point iteration and compute $\pi_{j}^{\text{new}}$ independently with simple vector operations that scale well for large data.
@@
2:$\bm{\theta}^{0}\leftarrow\underset{\bm{\theta}\in\Theta}{\arg\min}\,\sum_{i=1}^{n}\ell_{\bm{\theta}}(\bm{z}_{i})$triangleright $\pi^{0}=1$
3:for $k=1,2,\dots$ do
@@
5:     $\bm{\pi}^{k}\leftarrow$ fixed-point iterations (17) $\triangleright$ E-step
6:     $\bm{\theta}^{k}\leftarrow\underset{\bm{\theta}\in\Theta}{\arg\min}\,\sum_{i=1}^{n}\pi^{k}_{i}\ell_{\bm{\theta}}(\bm{z}_{i})$triangleright M-step
7:     if $\lVert\bm{\theta}^{k}-\bm{\theta}^{k-1}\rVert\leq tolerance$ then
@@
Stochastic approximation. The formulated objective $\mathcal{L}(\bm{\theta},\bm{\pi})$ alleviates the need to know or assess $\varepsilon$, which is particularly useful when $\varepsilon$ is not fixed, such as in the online learning setting. Another useful property of this variational bound is its structure – with respect to $\bm{\theta}$, the function consists of $n$ independent terms, just as the standard likelihood function, which can be employed by stochastic optimization of $\mathcal{L}(\bm{\theta},\bm{\pi})$. That is, if one does not have access to the full dataset but only to its batches, the M-step in Algorithm 1 can be replaced with a stochastic gradient descent (SGD) update,
@@
where the loss function and, therefore, the gradient components for each batch of $b$ samples $\{\bm{z}_{i}\}_{i=1}^{b}$ are re-weighted with $\bm{\pi}^{k}\in[0;1]^{b}$. The latter are computed using the corresponding loss values $\ell_{\bm{\theta}}(\bm{z}_{i}),\,\bm{z}_{i}=1,\dots,b$, at the preceding E-step with (17). Batch minimization of $\mathcal{L}(\bm{\theta},\bm{\pi})$, in effect, corresponds to a stochastic variant of the EM algorithm (see stepwise EM in [16]).
Truncation as a form of regularization. Since the objective in RLVI is suitable for stochastic optimization, our robust learning approach can also be used in deep learning, where SGD-like approaches are dominant. However, neural networks are overparameterized models and thus are prone to overfitting, which in theory (and in practice – see Figure 5) hinders the performance of RLVI. Indeed, if we substitute $\ell_{\bm{\theta}}(\bm{z}_{i})\approx 0$ for all samples into the stationary point condition (16), we find the corresponding minimum for $\bm{\pi}$,
$\pi^{\star}_{i}=\left(1+\frac{\varepsilon}{1-\varepsilon}e^{\ell_{\bm{\theta}}(\bm{z}_{i})}\right)^{-1}=1-\varepsilon,\quad i=1,\dots,n.$ (19)
That is, as soon as overfitting commences, the marginal likelihood approach treats corrupted samples as non-corrupted since overparameterized models are able to attain a high likelihood for outliers as well. Nevertheless, as is generally the case, to prevent overfitting, one can use regularization of the loss function. In this work, we introduce regularization to the RLVI algorithm based on the obtained posterior approximation, making the algorithm effective in the overparameterized regime as well. Namely, samples that have a low probability of being non-corrupt, $\pi_{i}<\tau$, are eliminated from SGD updates. This hard truncating approach raises two questions – when to apply it and how to select the decision boundary $\tau$. Both questions are formally resolved with no additional parameters required, as described in Section 4.3.

### v2

# Adaptive Robust Learning using Latent Bernoulli Variables

@@ Abstract
We present an adaptive approach for robust learning from corrupted training sets. We identify corrupted and non-corrupted samples with latent Bernoulli variables and thus formulate the learning problem as maximization of the likelihood where latent variables are marginalized. The resulting problem is solved via variational inference, using an efficient Expectation-Maximization based method. The proposed approach improves over the state-of-the-art by automatically inferring the corruption level, while adding minimal computational overhead. We demonstrate our robust learning method and its parameter-free nature on a wide variety of machine learning tasks including online learning and deep learning where it adapts to different levels of noise and maintains high prediction accuracy.
Machine Learning, ICML
@@ 1 Introduction
Several statistical learning problems are formulated as estimation of parameters $\bm{\theta}\in\Theta$ of a probabilistic model by maximizing its likelihood function $\prod_{i=1}^{n}p(\bm{z}_{i}|\bm{\theta})$ given $n$ independent observations $Z=\{\bm{z}_{i}\}_{i=1}^{n}$, $\bm{z}_{i}\sim p(\bm{z})$. By defining the loss function as the negative log-likelihood $\ell_{\bm{\theta}}(\bm{z})=-\ln p(\bm{z}|\bm{\theta})$, one can equivalently solve
@@
However, real-world data is often corrupted: samples arise from the true distribution $p(\bm{z})$ and a corrupting source $q(\bm{z})$,
$\bm{z}_{i}\sim(1-\varepsilon)p(\bm{z})+\varepsilon q(\bm{z}),$ (2)
@@
leading to a suboptimal solution $\bm{\theta}_{\text{ML}}$. In this contaminated mixture, corrupted samples $\bm{z}\sim q(\bm{z})$ may result from inaccurate measurements, errors or oversight in data acquisition or labeling, or even malicious attacks. Further, the corrupting distribution $q(\bm{z})$ is typically unknown, and the level of corruption $\varepsilon$, difficult to determine. Equation (2) used herein follows the well-known Huber contamination model from the robust statistics literature (Huber, 2011; Maronna et al., 2019).
Existing approaches require an estimation of the noise structure or the corruption level $\varepsilon$. This imposes restrictions on using these methods in settings where such pre-processing becomes impractical, e.g., in online learning, wherein the assumption about $\varepsilon$ needs to be optimized in continuously arriving data. In Figure 1, we show an example: the accuracy of binary classification trained on data that is continuously collected from the HAR dataset (Helou, 2023), and subsequently corrupted with varying number of randomly flipped labels in each batch (details in Section 4.2).
@@
Figure 1: Classification of online streaming data with varying number of corrupted labels. Adaptive nature of our approach (RLVI) allows for automatic identification of outliers when learning from batches of data with different $\varepsilon$. Our method is robust and thus has higher accuracy than the standard stochastic optimization of the likelihood (SGD).
Contribution. This paper presents a principled approach for robust learning from corrupted data that is
  • widely applicable given any likelihood function,
  • robust against a wide class of contamination sources,
  • adaptable and tuning-parameter free,
  • scalable for large data sets and deep learning models.
@@
Robust Learning via Variational Inference. The proposed approach, denoted as RLVI, employs latent Bernoulli variables to identify corrupted and non-corrupted training samples. The distribution of the latent variables is characterized from training data by maximizing the variational lower bound of marginal likelihood (variational inference) (Murphy, 2023). This allows detection of corrupted samples and computation of the optimal corruption level $\varepsilon$ automatically, subsequently learning the model $\bm{\theta}$ in a robust way. The problem formulation entailing RLVI is theoretically simple and also leads to an efficient and straightforward implementation.
Related work. While some approaches address learning from corrupted data with a special loss function, e.g., the Huber loss (Huber, 1992), others construct certain criteria to separate samples from the true distribution and the corrupted ones (Bhatia et al., 2015, 2017). Recently, new general approaches, such as Sever (Diakonikolas et al., 2019) and Robust Risk Minimization (RRM) (Osama et al., 2020), were introduced. Both methods are not model-specific. Sever is a general gradient-based learning method that penalizes the gradient components of the loss function corresponding to outliers. RRM involves obtaining sample weights from the constraint on the entropy of a weighted empirical distribution and minimizing the modified empirical risk. Robust Risk Minimization does not require calibrating multiple hyperparameters, unlike Sever. Despite not requiring the exact value of $\varepsilon$, both approaches are reliant on its upper-bound estimate $\widetilde{\varepsilon}\geq\varepsilon$. A very relevant work is (Wang et al., 2017), where the authors introduce latent variables $\bm{w}$ and raise likelihood terms to these variables to account for the departure from model’s assumptions. The paper suggests to impose a prior distribution for $\bm{w}$ that suits the problem at hand and infer $\bm{w}$ together with the model parameters. Thus, being quite general, this framework provides much freedom to users while leaving inference of $\bm{w}$ as a computationally challenging task solved with probabilistic programming. In this paper, we rather consider a specific case and define Bernoulli latent variables based on the Huber model (2), which enables efficient computation and does not require to specify any hyperparameters for a prior distribution.
In the recent past, learning from corrupted data has also received attention in the field of deep learning, and several approaches have been proposed to train a neural network in a robust way. For instance, one can use the correction of the regular loss function (Zhang & Sabuncu, 2018; Patrini et al., 2017). Alternatively, some approaches directly estimate the noise transition matrix (Goldberger & Ben-Reuven, 2016; Patrini et al., 2017). Others utilize an additional neural network to identify and prevent overfitting of the initial model, as in (Jiang et al., 2018), and to obtain a better accuracy by exchanging the information between two simultaneously trained networks, as in (Han et al., 2018) and (Wei et al., 2020). Approach from (Ren et al., 2018) is based on re-weighting the loss terms for each observation, wherein sample weights are treated as hyperparameters optimized with additional gradient computations. Another option is to combine techniques that are specific to deep learning: early-stopping (Xia et al., 2020) or dropout (Xu et al., 2023), with a criterion that aims to eliminate corrupted samples. However, training a neural network in the regular context where the contamination issue is not addressed, already requires certain hyperparameters to improve generalization, e.g., learning rate, batch size, number of layers, etc. The aforementioned approaches introduce additional parameters to combat the problem of corrupted labels, including $\varepsilon$, which makes their performance dependent on the efficacy of these parameters and the underlying assumptions.
@@ 2 Problem Formulation
Central to our approach is the introduction of a latent variable $t_{i}$ for each observation $\bm{z}_{i}\in Z$ such that
@@
Knowing the values of these variables will allow avoiding minimization of the losses on corrupted samples in the dataset in (1) by simply dropping the corresponding terms from the sum. In other words, latent variables enable us to define a likelihood function with respect to the non-corrupted data, such that
@@
where $\bm{t}=(t_{1},t_{2},\dots,t_{n})$. Optimization with respect to this likelihood function implies a combinatorial search over the latent variables $\bm{t}$, which is computationally infeasible even for moderate $n$.
@@
$\widehat{\bm{\theta}}=\underset{\bm{\theta}}{\arg\max}\,\max_{\varepsilon}p(Z|\bm{\theta},\varepsilon),$ (7)
addresses the problem of robust learning of $\bm{\theta}$ while obviating the need for specifying $\varepsilon$ or a combinatorial search over $\bm{t}$. We now turn to developing a method to approximately solve (7) in an efficient manner.
@@ 3 Method
We begin by introducing a variational bound on the marginal likelihood in (7) and then turn to implementation-specific aspects of the derived method.
Variational inference. To solve the formulated optimization problem, we need a tractable way to compute the sum over $\bm{t}$ in (6). Using Bayes’ rule, we express the marginal likelihood as
@@
where the denominator is the posterior distribution of latent variables $\bm{t}$ given the data $Z$, specified model $\bm{\theta}$, and the corruption level $\varepsilon$. As this posterior is intractable, we consider its variational approximation denoted $r(\bm{t}|\bm{\pi})$. Specifically, we use a distribution of $n$ independent Bernoulli variables with probabilities $\bm{\pi}=(\pi_{1},\dots,\pi_{n})$,
@@
Therefore, instead of maximizing the marginal likelihood function $p(Z|\bm{\theta},\varepsilon)$ directly, we apply the variational inference framework and optimize the so-called evidence lower-bound (ELBO) (Murphy, 2023). This lower bound has the following general form,
$\ln{p(Z|\bm{\theta},\varepsilon)}\geq\underbrace{\mathbb{E}_{r(\bm{t}|\bm{\pi})}\Big{[}\ln{p(Z|\bm{\theta},\bm{t})}\Big{]}-\text{KL}\Big{[}r(\bm{t}|\bm{\pi})\big{|}\big{|}p(\bm{t}|\varepsilon)\Big{]}}_{\text{ELBO}(\bm{\theta},\,\bm{\pi},\,\varepsilon)},$ (10)
@@
Note that corruption level does not appear in the first term of the ELBO and we can optimize $\varepsilon$ directly and independently of the model $\bm{\theta}$:
$\varepsilon=\arg\max\text{ELBO}(\bm{\theta},\bm{\pi},\varepsilon)=1-\frac{1}{n}\sum\limits_{i=1}^{n}\pi_{i}.$ (13)
That is, variational inference framework resolves an unknown hyperparameter by explicitly optimizing the corruption level. Moreover, we get an intuitively satisfying result that parameters $\pi_{i}$ defining the probability of each sample being non-corrupted, should sum up to the number of non-corrupted samples:
$n(1-\varepsilon)=\sum\limits_{i=1}^{n}\pi_{i}.$ (14)
Hence we arrive at the following objective – the negative ELBO, expressed as
$\mathcal{L}(\bm{\theta},\bm{\pi})=\sum\limits_{i=1}^{n}\pi_{i}\ell_{\bm{\theta}}(\bm{z}_{i})+\pi_{i}\ln\frac{\pi_{i}}{\langle\bm{\pi}\rangle}+(1-\pi_{i})\ln\frac{1-\pi_{i}}{1-\langle\bm{\pi}\rangle},$ (15)
where $\langle\bm{\pi}\rangle:=\sum_{i=1}^{n}\pi_{i}\,/\,n$ is the average of Bernoulli probabilities.
Consequently, the objective in (7) is replaced by the optimum of its variational bound, i.e.,
$\boxed{\bm{\theta}_{\text{RLVI}}=\underset{\bm{\theta}\in\Theta}{\arg\min}\,\min_{\bm{\pi}\in(0;1)^{n}}{\mathcal{L}(\bm{\theta},\bm{\pi})}.}$ (16)
Numerical optimization. The resulting optimization problem is solved using the block-wise Algorithm 1, which follows the general Expectation-Maximization (EM) scheme (Bishop & Nasrabadi, 2006). The E-step consists of minimizing $\mathcal{L}(\bm{\theta},\bm{\pi})$ in $\bm{\pi}$ for fixed parameters $\bm{\theta}$. In the M-step, the inferred probabilities are used to maximize the re-scaled log-likelihood, $-\sum_{i=1}^{n}\pi_{i}\ell_{\bm{\theta}}(\bm{z}_{i})$. As an example, in case of linear regression the latter step solves a weighted least-squares problem. For a classification task, we minimize the cross-entropy loss corrected with sample weights.
Note that the E-step can be performed efficiently, as the objective $\mathcal{L}(\bm{\theta},\bm{\pi})$ is convex in $\bm{\pi}$. Therefore, to find the optimal parameters $\bm{\pi}$ for a fixed model, the derivative of $\mathcal{L}(\bm{\theta},\bm{\pi})$ is equated to zero w.r.t. $\pi_{j}$ for all $j=1,\dots,n$,
@@
This is a system of nonlinear equations for $\bm{\pi}$, which we solve with the fixed-point iterations,
$\pi_{j}^{\text{new}}=\left(1+\dfrac{1-\langle\bm{\pi}\rangle^{\text{old}}}{\langle\bm{\pi}\rangle^{\text{old}}}e^{\ell_{\bm{\theta}}(\bm{z}_{j})}\right)^{-1}.$ (18)
In (18), for the mean weight, we use the value from the previous fixed-point iteration and compute $\pi_{j}^{\text{new}}$ independently with simple vector operations that scale well for large data. A proof for $\mathcal{L}(\bm{\theta},\bm{\pi})$ being convex in $\bm{\pi}$ and iterations (18) converging to a stationary point is provided in the Appendix.
@@
2:  $\bm{\theta}^{0}\leftarrow\underset{\bm{\theta}\in\Theta}{\arg\min}\,\sum_{i=1}^{n}\ell_{\bm{\theta}}(\bm{z}_{i})$ // $\pi^{0}_{i}=1$
3:  for $k=1,2,\dots$
@@
5:     $\bm{\pi}^{k}\leftarrow$ fixed-point iterations (18) // E-step
6:     $\bm{\theta}^{k}\leftarrow\underset{\bm{\theta}\in\Theta}{\arg\min}\,\sum_{i=1}^{n}\pi^{k}_{i}\ell_{\bm{\theta}}(\bm{z}_{i})$ // M-step
7:     if $\lVert\bm{\theta}^{k}-\bm{\theta}^{k-1}\rVert\leq tolerance$
@@
9:     end if
10:  end for
Stochastic approximation. The formulated objective $\mathcal{L}(\bm{\theta},\bm{\pi})$ alleviates the need to know or assess $\varepsilon$, which is particularly useful when $\varepsilon$ is not fixed, such as in the online learning setting. Another useful property of this variational bound is its structure: with respect to $\bm{\theta}$, the function consists of $n$ independent terms, just as the standard likelihood function, which can be employed by stochastic optimization of $\mathcal{L}(\bm{\theta},\bm{\pi})$. If one does not have access to the full dataset but only to its batches, the M-step in Algorithm 1 can be replaced with a stochastic gradient descent (SGD) update
@@
where the loss function and, therefore, the gradient components for each batch of $b$ samples $\bm{z}_{i}$ are re-weighted with $\pi^{k}_{i},\,i=1,\dots,b$. The latter are computed using the corresponding loss values $\ell_{\bm{\theta}}(\bm{z}_{i})$ at the preceding E-step with (18). Batch minimization of $\mathcal{L}(\bm{\theta},\bm{\pi})$, in effect, corresponds to a stochastic variant of the EM algorithm – see, e.g., stepwise EM in (Murphy, 2023).
Truncation as a form of regularization. Since the objective in RLVI is suitable for stochastic optimization, our robust learning approach can also be used in deep learning, where SGD-like approaches are dominant. However, neural networks are overparameterized models and thus are prone to overfitting, which in theory (and in practice: see Figure 5) hinders the performance of RLVI. Indeed, if we substitute zero loss for all training samples into the stationary point condition (17), we find the corresponding minimum for all $i=1,\dots,n$,
$\pi^{\star}_{i}=\left(1+\frac{\varepsilon}{1-\varepsilon}e^{\ell_{\bm{\theta}}(\bm{z}_{i})}\right)^{-1}\Bigg{|}_{\ell_{\bm{\theta}}(\bm{z}_{i})=0}=1-\varepsilon.$ (20)
That is, when overfitting commences, the marginal likelihood approach treats all samples as non-corrupted since overparameterized model is capable of minimizing loss to zero on both ‘clean’ and corrupted samples. Nevertheless, as is generally the case, to prevent overfitting, one can use regularization of the loss function. In this work, we introduce regularization to the RLVI algorithm, making the algorithm effective in the overparameterized regime as well. Namely, samples that have a low probability of being non-corrupt are eliminated from SGD updates:
$\pi_{i}<\tau\implies\pi_{i}\leftarrow 0.$ (21)
Furthermore, we define the threshold $\tau$ based on the following criterion: maximize the number of samples to be used for learning (maximize $\tau$) subject to a bounded type II error (number of corrupted samples treated as ‘clean’). Hence, $\tau=\max\,\{\pi_{1},\pi_{2},\dots,\pi_{n}\}$, such that
$\frac{\mathbb{E}_{r}\Big{[}\#\text{False Clean}\Big{]}}{\mathbb{E}_{r}\Big{[}\#\text{Corrupted \Big{]} }}=\frac{\sum\limits_{i=1}^{n}(1-\pi_{i})\mathbbm{1}[\pi_{i}\geq\tau]}{\sum\limits_{i=1}^{n}(1-\pi_{i})}\leq 0.05.$ (22)
In this criterion, the admissable type II error is common for statistical hypothesis testing and equals 5%. Also note how the obtained posterior approximation is employed: the numerator is the expected number of corrupted samples considered as ‘clean’, and the denominator is the expected total number of corrupted samples, based on $r(\bm{t}|\bm{\pi})$.
The resulting variant of RLVI, to be used for overparameterized models, is listed as Algorithm 2. It implements RLVI as stochastic gradient optimization of neural network parameters $\bm{\theta}$. Parameters $\pi_{i}$ are updated at the end of each epoch using efficient iterations (18). To prevent model’s overfitting to corrupted samples, the algorithm eliminates gradient terms corresponding to low $\pi_{i}$. Threshold for truncation is re-computed across epochs as a non-decreasing value from the type II error criterion (22). Note that Bernoulli probabilites are updated once at each epoch, which makes solution for $\bm{\pi}$ less dependent on the batch size in this case.